Queenb7282:

juliana lives and works in pennsylvania which has a flat state income of 3.07%. if her annual salary is $41,995 and if she gets paid once a month how much is withheld from her gross income for state income tax each pay period?

justus:

Her total annual income is 41,995

justus:

If we want to get her monthly income we have to do her annual income over 12

justus:

So it would be 41,995 over 12

justus:

Her tax rate per month = 3.07%

justus:

\[ 3.07\%\text{ of }\frac{41995}{12}\]\]

justus:

\[ \frac{3.07}{100}\times \frac{41995}{12}\]

justus:

\[ \frac{128924.65}{1200}\]

justus:

\[ = 107.437208\approx 107.44\]